if sec∂ and cosec∂ are the roots of x2-px+q=0 then
  1. p2=q(q-2)
  2. p2+q2=2q
  3. p2=q(q+1)
  4. p2=q(q-1)
  5. p2=q(q+2)       pls explain

Grade:12th Pass

2 Answers

ASh
6 years ago
guys please answer................................................................................................................
 
 
amit kumar
6 years ago
Sum of roots seca+coseca=p.....1And product seca.coseca=q......2From 2 sina.cosa=1/qFrom 1 sina+cosa/sina.cosa=sina+cosa)q=pSquiringSin²a+cos²a+2sina.cosa=p²/q²So 1+2/q=p²/q²Which on simplification givesP²=q²+2q=q(q+2)Last choice